# Unraveling the Smoothness Properties of Diffusion Models: A Gaussian Mixture Perspective

### v1

@@ Appendix I More Calculation for Application
Now we want to find a more concrete $L$ in Assumption D.1. Notice that from Lemma H.1, we know that at any time between $0\leq t\leq T$, $p_{t}$ is also a $k$-mixture of gaussian, except that the mean and covariance change with time.
Using Lemma H.3, we can get $L$.
Now we want to find the second momentum in Assumption D.2. Using Lemma H.2, we know that $m_{2}=(\sum_{i=1}^{k}\alpha_{i}(\|\mu_{i}\|_{2}^{2}+\operatorname{tr}[\Sigma_{i}]))^{1/2}$.
@@
Now we want to have a more concrete setting for Theorem D.5 by calculating each term directly. Notice that now we have all the quantities except for the KL divergence term. Thus, we calculate $\mathrm{KL}(p_{0}\|\mathcal{N}(0,I))$, which means the KL divergence of data distribution and standard Gaussian.
@@
Then we have all the quantities in Theorem D.5. After directly applying the theorem, we finish the proof. ∎
I.2 Concrete calculation for Theorem D.6
Theorem I.2 ($\mathsf{DDPM}$, KL divergence, formal version of Theorem 5.7).
@@
Using Lemma H.3, we can get $L$. Using Lemma H.2, we can get $m_{2}$. Then we directly apply Theorem D.6. ∎
I.3 Concrete calculation for Theorem D.7
Theorem I.3 ($\mathsf{DDPM}$, KL divergence for smooth data distribution, formal version of Theorem 5.8).
@@
Clearly, $p_{0}$ is second-order differentiable. Using Lemma H.3, we can get $L$. Using Lemma H.2, we can get $m_{2}$. Then we directly apply Theorem D.7. ∎
I.4 Concrete calculation for Theorem D.8
Theorem I.4 ($\mathsf{DPOM}$, formal version of Theorem 5.9).
@@
Using Lemma H.3, we can get $L$. Using Lemma H.2, we can get $m_{2}$. Then we directly apply Theorem D.8. ∎
I.5 Concrete calculation for Theorem D.9
Theorem I.5 ($\mathsf{DPUM}$, formal version of Theorem 5.10).
@@
Using Lemma H.3, we can get $L$. Using Lemma H.2, we can get $m_{2}$. Then we directly apply Theorem D.9. ∎

### v2

@@ Appendix H More Calculation for Application
Now we want to find a more concrete $L$ in Assumption 6.1. Notice that from Lemma G.1, we know that at any time between $0\leq t\leq T$, $p_{t}$ is also a $k$-mixture of gaussian, except that the mean and covariance change with time.
Using Lemma G.3, we can get $L$.
Now we want to find the second momentum in Assumption 6.2. Using Lemma G.2, we know that $m_{2}=(\sum_{i=1}^{k}\alpha_{i}(\|\mu_{i}\|_{2}^{2}+\operatorname{tr}[\Sigma_{i}]))^{1/2}$.
@@
Now we want to have a more concrete setting for Theorem 6.5 by calculating each term directly. Notice that now we have all the quantities except for the KL divergence term. Thus, we calculate $\mathrm{KL}(p_{0}\|\mathcal{N}(0,I))$, which means the KL divergence of data distribution and standard Gaussian.
@@
Then we have all the quantities in Theorem 6.5. After directly applying the theorem, we finish the proof. ∎
H.2 Concrete calculation for Theorem 6.6
Theorem H.2 ($\mathsf{DDPM}$, KL divergence, formal version of Theorem 5.7).
@@
Using Lemma G.3, we can get $L$. Using Lemma G.2, we can get $m_{2}$. Then we directly apply Theorem 6.6. ∎
H.3 Concrete calculation for Theorem 6.7
Theorem H.3 ($\mathsf{DDPM}$, KL divergence for smooth data distribution, formal version of Theorem 5.8).
@@
Clearly, $p_{0}$ is second-order differentiable. Using Lemma G.3, we can get $L$. Using Lemma G.2, we can get $m_{2}$. Then we directly apply Theorem 6.7. ∎
H.4 Concrete calculation for Theorem 6.8
Theorem H.4 ($\mathsf{DPOM}$, formal version of Theorem 5.9).
@@
Using Lemma G.3, we can get $L$. Using Lemma G.2, we can get $m_{2}$. Then we directly apply Theorem 6.8. ∎
H.5 Concrete calculation for Theorem 6.9
Theorem H.5 ($\mathsf{DPUM}$, formal version of Theorem 5.10).
@@
Using Lemma G.3, we can get $L$. Using Lemma G.2, we can get $m_{2}$. Then we directly apply Theorem 6.9. ∎